(20 points) Apply 3 iterations by hand (but use a regular calculator) of the multi-variable Newton method to estimate the root of the following equation system

x^ 2 + y^ 2 = 3

xy = 1
Answer #1

Answer:

x² + y² -3=0 xy-1=0 fi fa 56) - LR ters] = 241 As per Hewtons multivariate method Xitt= xi - (J(xi)] [ f(x;)] Iteration 1: XIteration 2 o [X2] = 1.665 - 3.33 3.334 | 2.5511 Y J 1.667 -1.667 1665 11-7756 | = 1.6657 - 10.15 1.50220550 L 1.667 - 10.15